OpenStudy (anonymous):

Solve the following systems of equations using the graphing method. What type of system is it? Name the solution if there is one. 2x-2y=8 x-y=4 show graph

OpenStudy (anonymous):

here is my work 2x-2y=8 -2x -2x -2y=-2x+8 /-2 /-2 y=x-4 x-y=4 add y: x = y + 4 subtract 4: x - 4 = y or y = x - 4

OpenStudy (anonymous):

i need help with the graph. what should graph look like?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@sammie2017

Nnesha (nnesha):

ok so they said using graphing method its mean solve for y and put value into the calculator where they said to graph ??

OpenStudy (anonymous):

right i have to solve for y to get into slope intercept form right..? @nnesha

OpenStudy (anonymous):

@aryandecoolest @sidsiddhartha @Dark_Riddles @fateal @Harsha19111999 @kohai

OpenStudy (fateal):

this is one of the graphs ;o

OpenStudy (anonymous):

@fateal is there suppose to be another line?

OpenStudy (fateal):

tbh yes, im not sure where tho :/ http://www.mathpapa.com/algebra-calculator.html check this web

OpenStudy (anonymous):

@zepdrix @xstehzy @Compassionate @vishweshshrimali5 @BassCatcher15 @nikato @m

OpenStudy (anonymous):

hmmm okay i wil check it out.. i need help i always have issue with graph. need answer for tomorrow.

OpenStudy (fateal):

this web will help you graph anything and everything :)

zepdrix (zepdrix):

What do you need help with spence? :d looks like you've figured it out for the most part. You determined that they're both the same line, one is just a multiple of the other, which means that they share infinitely many points, every point as a solution.

OpenStudy (anonymous):

i think it is consistent since they are the same answer?

OpenStudy (anonymous):

or would that be infinite

OpenStudy (anonymous):

@fateal

OpenStudy (fateal):

its really late here, ive had a full day, tbh i cant focus, i was about to leave but you tagged me, and i think it will be infinte since if you noticed the fact that both of them have the same points

OpenStudy (fateal):

night

OpenStudy (anonymous):

okay i appreeciate help get some sleep!